If x+1/x ≠ 0

x^3 + 1/ x^3 =0

find x^5+ 1/ x^5
1
Expert's answer
2013-01-14T09:04:40-0500
x^3 + 1/ x^3 =0
x^3 = -1/x^3
x=-1/x

hence

x^5+1/x^5 = x^5-x^5 = 0